26.4 The frangle industry is a monopoly, with a demand curve 100-p, where p is the price of

frangles. It takes one unit of labor and no other inputs to produce a frangle. The Frangle-makers

Guild is a strong union. The Guild sets a wage and prevents anyone from working for less than

that wage. The frangle monopoly must pay that wage but can hire as much labor as it chooses to.

If the guild chooses a wage so as to maximize the total earnings (wage times number of units of

labor hired) of frangle-makers, then:

(a) the price of frangles will be 50.

(b) the price of frangles will be 25.

(c) the price of frangles will equal the wage rate.

(d) the wage rate will be 25.

(e) the wage will be 50.

The answer is e, why?
Answer #1

Answer is option E)

now maximize wL

Then , labor demand curve :

MRPL = w

MR*MPL = w

As, P = 100-Q

MR = 100-2Q

Q = L : Production function

MPL = 1

So, 100-2L = w

So, z = total labor earning = wL

z = (100-2L)*L

z = 100L - 2L2

FOC:

dz/dL = 100-4L = 0

so L* = 25

w* = 100-2*25 = 50

so P* = 100-25 = 75